OpenStudy (anonymous):

2r^2-14r+24 answer please

Directrix (directrix):

3 and 4

Directrix (directrix):

2r^2 - 14 r + 24 =0 r^2 - 7r + 12 = 0 r -3 = 0 or r-4 = 0 r = 3 or r = 4

Directrix (directrix):

Factor: 2(r-3)(r-4)

Directrix (directrix):

Chiamak --> It would be good to include instructions with the problem. Do you want to factor or what? Thanks.